# SpargeAttention: Accurate and Training-free Sparse Attention Accelerating Any Model Inference

### v5

@@ 3.1 Sparse FlashAttention
where $m_{ij}$ and $l_{ij}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an operator similar to softmax.: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$. Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{ij})^{-1}O_{ij}$.
@@
15:          if $\max(m_{\text{local}}[I_{w}]-m_{ij}[I_{w}])<\lambda$  then
@@ 3.2 Selective Token Compression for Sparse Prediction
Prediction. As shown in Step1 in Fig. 3, we first compute a mean cosine similarity across tokens for each block of $Q$ and $K$. Next, we compress each block into a single token by calculating a mean across tokens. Then, we compute a compressed $QK^{\top}$ using the compressed $Q$ and $K$. Finally, to prevent interference from non-self-similar blocks, i.e., the block similarity less than a hyper-parameter $\theta$, we set the corresponding values in $S$ to $\infty$- ∞, and then obtain a compressed attention map through softmax. This algorithm can be expressed as:
@@ 3.7 HilbertCurve Permutation
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ WikiText (Ppl.) $\downarrow$ Longbench $\uparrow$ InfiniteBench $\uparrow$ NIAH $\uparrow$
@@
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
@@
        Model (seq_len)         Attention (Sparsity)        Speed (TOPS)$\uparrow$        FID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
@@ 4.3 Ablation Study and key Insights
Strategy only $M_{g}$ only $M_{pv}$ $M_{g}$ +$M_{pv}$
@@
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$+$M_{pv}$ on Llama3.1 in Needle-in-a-Haystack task with 128K sequence length.
@@ A.2 Ablation Study of Self-Similarity Judge
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ NIAH $\uparrow$

### v6

@@ 3.1 Sparse FlashAttention
[…]
@@
15:    if $\max(m_{\text{local}}[I_{w}]-m_{ij}[I_{w}])>\lambda$  then
@@ 3.2 Selective Token Compression for Sparse Prediction
Prediction. As shown in Step1 in Fig. 3, we first compute a mean cosine similarity across tokens for each block of $Q$ and $K$. Next, we compress each block into a single token by calculating a mean across tokens. Then, we compute a compressed $QK^{\top}$ using the compressed $Q$ and $K$. Finally, to prevent interference from non-self-similar blocks, i.e., the block similarity less than a hyper-parameter $\theta$, we set the corresponding values in $S$ to -$\infty$, and then obtain a compressed attention map through softmax. This algorithm can be expressed as:
@@ 3.7 HilbertCurve Permutation
[…]
@@
[…]
@@
[…]
@@ 4.3 Ablation Study and key Insights
Strategy only $M_{g}$ only $M_{pv}$ $M_{g}$ $M_{pv}$
@@
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$$M_{pv}$ on Llama3.1 in Needle-in-a-Haystack task with 128K sequence length.
@@ A.2 Ablation Study of Self-Similarity Judge
[…]